Mr.baker creates 35 math test problems in 5 hours for the upcoming math pop quiz.How many questions does Mr.baker create in 3 hours?
Answer by math_helper(2461)


———
35 questions/5hrs = 7 questions/hr
7 questions/hr * 3hrs = 21 questions.
———
Alternatively: x/3 = 35/5 —> x = 35*3/5 = 21